OpenStudy (anonymous):

3x - y = - 4 x + y = 8 For these 2 problems what would be the most convenient method to use? A. Graphing B. Substitution C. Elimination

OpenStudy (anonymous):

ELIMINATION

OpenStudy (anonymous):

ANSWER IS X= 1 AND Y= 7

OpenStudy (anonymous):

3x-y=-4 x+y=8 so 4x= 4 x=1 3+y=8 so y=5 so answer is X=1& Y=5

OpenStudy (anonymous):

y is not equal to 5 as 3*1-5 is not equal to -4.

OpenStudy (anonymous):

oh yes 1+y=8 not 3 thanx

OpenStudy (anonymous):

hmm.